(1) At which angle o does the polar curve r = 1 + 2 sin 0 intersect the origin? 117 (2) Draw the polar curve r=1+ 2 sin 8. LeThe answers in the box are not necessarily correct.
Answer #1


BB r=1+2 SING for intersection at origin 1+2Sin 0 =0 Sino=-112 = Sin (7%) oznatin(78%) 0=72,144. - -- A=2 22.do